What is the absolute value of 5-3 and 4-|4-2|?
Answer by jim_thompson5910(35256)

|5-3| = |2| = 2


So |5-3| = 2

-------------------------------------------------------

4-|4-2| = 4 - |2| = 4 - 2 = 2


So 4-|4-2| = 2